Bill cornett won $200,000 in a state lottery. He first paid income tax of 30% on the winnings. Of the rest, he invested some at 8.5% and some at 7%, making $10,700 interest per year. How much was invested at each rate?
Answer by mathstutor458(57)

Bill cornett amount won in lottery = $200,000
He paid income tax of 30% on the winnings.
Therefore, amount paid=30% of 200,000=$60,000
Remaining Amount=$200,000-$60,000=$140,000
he invested some,let us say x at 8.5% . amount invested at 8.5%= 8.5*x/100
he invested some,let us say x at 7% . amount invested at 7%= 7*(140000-x)/100
At the end of the year, he got $10,700 interest.
therefore,8.5x/100+7(140000-x)/100=10,700
By solving this equation
we get,x=$60,000 which is the amount invested at 8.5%
140000-60000=$80000 is the amount invested at 7%